# Towards Faithful and Robust LLM Specialists for Evidence-Based Question-Answering

### 16 Feb 2024 v2

@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023; Anil et al., 2023) have become the center of many cutting-edge applications due to their generalisability and information processing abilities. A typical application of LLMs is in Evidence-Based Question Answering (QA), where LLMs are expected to answer questions based on provided sources and cite the sources accurately (e.g., Ni et al., 2023; Vaghefi et al., 2023; Cui et al., 2023; Liu et al., 2023a). By providing these additional sources, multiple shortcomings of standalone LLMs, such as hallucination (Ji et al., 2023) and limited knowledge capacity (Hu et al., 2023), can be addressed, thereby enhancing answer traceability (Gao et al., 2024).
However, the performance of existing LLMs on Evidence-Based QA is far from perfect. The SOTA close-sourced LLMs and generative search engines have an unignorable rate of hallucinated answers and false citation (Ni et al., 2023; Liu et al., 2023a). Unfortunately, open-sourced LLMs are even less faithful than the already quality-lacking close-sourced LLMs in Evidence-Based QA (Yue et al., 2023; Gao et al., 2023; also see our evaluation in Section 5.1), although they achieve competitive results on general instruction-following benchmarks (Touvron et al., 2023; Tunstall et al., 2023). We argue that this may prevent practitioners from building Evidence-Based QA (or other RAG) applications in a robust way. Therefore, efficient data creation and fine-tuning methods are urgently needed for practitioners to improve LLMs’ Evidence-Based QA performance in their target applications.
@@
To address C1, we propose a data generation pipeline that synthesizes SynSciQA (Synthetic Scientific Question Answering), a well-diversified synthetic dataset for Evidence-Based QA. Then we extend the pipeline with two novel quality filters to sift out low-quality synthetic data points, leading to SynSciQA+ and SynSciQA++ (see the left half of Figure 1). To address C2, we first collect an in-domain test set SynSciQA with the data generation pipeline, which shares the data distribution with the training data (i.e., SynSciQA) but covers different topics. We further collect three test sets with different distances to the training data distribution to study the OOD performance (see the right half of Figure 1).
Extensive experiments on all proposed train and test settings show that (1) data quality is more important than quantity in Evidence-Based QA fine-tuning; (2) fine-tuning on generated data improves the performance on both in- and out-of-distribution test sets; and (3) performance scores on in-domain test set substantially indicate the OOD performance, suggesting that practitioners may use the synthetic data for validation to estimate the OOD performance. All evaluation metrics are based on golden heuristics and best-performed models from previous work (Yue et al., 2023), which we further verified with human and GPT-4 evaluation. In summary, our contributions include:
@@ 2.2 Quality Dimensions
Our quality dimensions and fine-tuning focus on faithfulness, the most significant shortcoming of open-sourced LLMs on Evidence-Based QA (Yue et al., 2023; Gao et al., 2023). Another important dimension is helpfulness, which can be defined as “how well does the answer address the question?”. Our quality dimensions partially address helpfulness by measuring truthful responses based on question-relevant sources. However, we argue that helpfulness is hard to define and evaluate objectively. For this task, it is also challenging to disentangle helpfulness from faithfulness, as a response can only have high helpfulness if it follows the prompt well, i.e., obeying all the faithful citation requirements. To shed light on this aspect, we put additional analyses in Appendix A.
@@ 2.3 Evaluation Metrics
where () denotes the collection of factually entailed (unentailed) sentences, and denotes the collection of answer sentences with a wrong format or without citation. While the format quality is easy to measure through heuristics, the answer’s sentence-source entailment is challenging and requires neural model prediction. In this work, we aggregate the best-performing attribution-prediction models of previous work: attrscore-flan-t5-xl and -xxl checkpoints from Yue et al. (2023) to measure entailment. To achieve higher precision, a sentence is entailed by the cited source only if both models predict “attributable”. The attributability score is not applicable for answers without any citation () since models should not cite when there is no relevant source. Those answers are addressed by source quality scores (i.e., the model should cite when there is a relevant citation). We mostly follow “citation recall”, the central metric of Gao et al. (2023), to design attributability scores but adjust it to our setting of Evidence-Based QA (more details in Appendix B).
@@ 3.2 Automated Quality Filters
The clear task definition and data creation process allow us to apply quality filters on the dataset. First, we apply a source quality filter to the original SynSciQA. Through its construction process, we know which sources are relevant and irrelevant to the question. Thus, we filter out data points that do not pass the source quality scores. This leaves us with 1386 samples which we call SynSciQA+.
@@ 4 Evaluation Datasets
Finally, we use another RAG tool to create ClimateQA. ClimateQA 333https://huggingface.co/spaces/Ekimetrics/climate-question-answering is a RAG system that answers questions based on IPCC and IPBES reports. We pose 261 climate-related questions from Welch (2022) to the system and store the outputted sources. Again, we use this data as input for our instruction form.
@@ 5.1 Zero-Shot Performance
We first use the proposed test sets and evaluation metrics to benchmark the zero-shot performance of close-sourced and open-sourced LLMs on Evidence-Based QA. The results are shown in Table 1. We find that there is a significant performance gap between open- and close-sourced LLMs on Evidence-Based QA, although they achieve comparable performance on general instruction-following benchmarks (for instance, Zephyr-7b- vs. GPT-3.5 on MT-Bench (Tunstall et al., 2023)). Similarly, Tunstall et al. (2023) shows that Zephyr-7b- outperforms Llama-2-70b-chat on all dimensions of MT-Bench, while our evaluation shows that Llama-2-13b-chat is less hallucinated than Zephyr-7b- on Evidence-Based QA. Therefore, the proposed Evidence-Based QA benchmarks can be an effective resource to benchmark LLMs’ faithfulness, supplementing MT-Bench.
@@ 5.2 SynSciQA Fine-Tuning
RQ2.2: Synthetic data as validation set for OOD performance. We observe a fluctuating performance corresponding to fine-tuning epochs. Therefore, it is important to conduct checkpoint selection over epochs with a validation (or development) set during fine-tuning. However, performance on SynSciQA is much higher than that of other OOD test sets. Therefore, the performance on a synthetic dataset cannot directly reflect the OOD or real-world performance. But can in-domain synthetic data still be an effective development set indicating which epoch may perform best on OOD data? To answer this question, we compute the Pearson’s Correlation between performance scores of different test sets. Results are presented in Table 3, illustrating that the performance on synthetic data has a strong correlation with OOD performance. However, the correlation becomes weaker when the distribution is more distant (ClimateQA and ChatReport has a weaker correlation than SynSciQA). Therefore, we conclude that synthetic data can provide a valid development set for OOD performance.
@@
RQ2.3: Overfitting does exist. We suspect that fine-tuning too many epochs may cause overfitting and reduce generalisability. So we compute the Pearson’s Correlation Coefficient between the performance scores and epoch numbers of all settings, where positive correlations indicate benefit from more epochs and negative correlations indicate overfits. Results are visualized in Figure 8, showing an overfitting trend for the majority of settings. Therefore, fine-tuning too many epochs may lead to a suboptimal performance. But we do not observe the fine-tuning overfits more to SynSciQA than others. We attribute this to SynSciQA containing different scientific topics from the training data.
@@ 5.3 Validating Attributability Score
Then, we calculate the Pearson correlations between NLI-model-based attributability scores and human / GPT-4 evaluated attributability. As Table 4 shows, the results substantiate the validity of our method for calculating answer attributability. Correlations exceeding 80% across all comparisons between our scores and those annotated by humans and GPT-4 affirm the mutual reinforcement of the outcomes. For more details, see Appendix I.
We also notice a potential shortcut to improve the attributability score: a model may only improve its format quality (i.e., providing a citation to more sentences or more correctly writing source names) without improving the answer entailment rate. In Appendix J, we provide the entailment ratio of format-correct citations as a side result to show that this shortcut does not exist. Both format and attributability dimensions are improved by fine-tuning.
| 0.821 | 0.917 | 0.871 |
@@ 6 Related Work
Basing Answers on Sources: Prompting LLMs to respond with citations has been a popular pattern of Retrieval Augmented Generation (RAG) for better traceability (Karpukhin et al., 2020; Lewis et al., 2021; Borgeaud et al., 2021; Vaghefi et al., 2023; Ni et al., 2023; Asai et al., 2023; Li et al., 2023; Gao et al., 2024). However, previous work shows that asking for citations does not make the answer more factually trustworthy (Min et al., 2023). Commercial search engines and SOTA closed-sourced LLMs suffer from unsatisfactory performance (Ni et al., 2023; Liu et al., 2023a), while open-sourced LLMs have even worse faithfulness (Yue et al., 2023; Gao et al., 2023). Therefore, LLMs, especially open-sourced ones, need essential improvement to achieve more trustworthy RAG applications.
@@ 7 Conclusion
Second, due to these budget and time limitations, we also conduct random sampling when performing human and GPT-4 evaluation to verify the attributability score instead of evaluating all instruction-answer pairs in all settings and epochs. However, we argue that the performance of different settings is uniform across different instructions and sampled data points are representative enough. Furthermore, we make all hand evaluations and generations of all checkpoints publicly available (for more details, see Appendix I).
@@
Data Privacy or Bias: There are no data privacy issues or bias against certain demographics with regard to the data collected from real-world applications and LLM generations.
@@ Appendix D Hand-Evaluation of the Quality Filters
On the other hand, 4% of the cases were split decisions. These predominately originate from different interpretations of nuances in the used language. Figure 10 shows an example of this.
@@ Appendix F Test Data Examples
The first evident difference lies in the statistical properties of the sources in the instructions (see Table 10). We create SynSciQA and GenSearch with the data creation pipeline described in section 2. This means, we have known relevant and irrelevant sources in the datasets. On the other hand, we use top-10 retrieved sources for ChatReport and the output sources by ClimateQA for ClimateQA. As Table 10 shows, the different sourcing mechanisms result in a large difference in average length and standard deviation.
@@
Appendix G Hyperparameters
@@
All experiments are conducted on two clusters, one with 4 V100 GPUs and the other with 4 A100 (80G) GPUs.
Appendix H Statistical Significance Tests
@@
Appendix I Validating the Attributability Score
The target of this validation is to reinforce the validity of the employed methodology in the answer attributability. Generally, this investigation follows the same structure as our train set validation in Appendix D. The main difference is that we now investigate all settings, including those out-of-distribution to find evidence that the comparisons are valid. To investigate the decisions, we repeat the evaluation of our score with both human and GPT-4 annotation. Both evaluations follow the structure of answer attributability and are articulated through the following prompt.
@@ Appendix I Validating the Attributability Score
Appendix J Format Short-Cut in Attributability

### 26 Feb 2024 v3

@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023; Anil et al., 2023) have become the center of many cutting-edge applications due to their generalisability and information processing abilities. A typical application of LLMs is Evidence-Based Question Answering (QA), where LLMs are expected to answer questions based on provided sources and cite the sources accurately (e.g., Ni et al., 2023; Vaghefi et al., 2023; Cui et al., 2023; Liu et al., 2023a). By providing these additional sources, multiple shortcomings of standalone LLMs, such as hallucination (Ji et al., 2023) and limited knowledge capacity (Hu et al., 2023), can be addressed, thereby enhancing answer traceability (Gao et al., 2024).
However, the performance of existing LLMs on Evidence-Based QA is far from perfect. The SOTA close-sourced LLMs and generative search engines have an unignorable rate of hallucinated answers and false citation (Ni et al., 2023; Liu et al., 2023a). Unfortunately, open-sourced LLMs are even less faithful than the already quality-lacking close-sourced LLMs in Evidence-Based QA (Yue et al., 2023; Gao et al., 2023; also see our evaluation in Section 5.1), although they achieve competitive results on general instruction-following benchmarks (Touvron et al., 2023; Tunstall et al., 2023). We argue that this may prevent practitioners from building Evidence-Based QA (or other RAG) applications in a robust way. Therefore, efficient data creation and fine-tuning methods are urgently needed to improve LLMs’ Evidence-Based QA performance in target applications.
@@
To address C1, we propose a data generation pipeline that synthesizes SynSciQA (Synthetic Scientific Question Answering), a well-diversified synthetic dataset for Evidence-Based QA. We further extend the pipeline with two novel quality filters to sift out low-quality synthetic data points, leading to SynSciQA+ and SynSciQA++ (see the left half of Figure 1). To address C2, we first collect an in-domain test set SynSciQA with the data generation pipeline, which shares the data distribution with the training data (i.e., SynSciQA) but covers different topics. We further collect three test sets with different distances to the training data distribution to study the OOD performance (see the right half of Figure 1).
Extensive experiments on all proposed train and test settings show that (1) data quality is more important than quantity in Evidence-Based QA fine-tuning; (2) fine-tuning on generated data improves the performance on both in- and out-of-distribution test sets; and (3) performance scores on in-domain test set substantially indicate the OOD performance, suggesting that the synthetic data can be used for validation to estimate the OOD performance. All evaluation metrics are based on golden heuristics and best-performed models from previous work (Yue et al., 2023), which we further verified with human and GPT-4 evaluation. In summary, our contributions include:
@@ 2.2 Quality Dimensions
Our quality dimensions and fine-tuning focus on faithfulness, the most significant shortcoming of open-sourced LLMs in Evidence-Based QA (Yue et al., 2023; Gao et al., 2023). Another important dimension is helpfulness, which can be defined as “how well does the answer address the question?”. Our quality dimensions partially address helpfulness by measuring truthful responses based on question-relevant sources. However, we argue that helpfulness is hard to define and evaluate objectively. For this task, it is also challenging to disentangle helpfulness from faithfulness, as a response can only have high helpfulness if it follows the prompt well, i.e., obeying all the faithful citation requirements. To shed light on this aspect, we put additional analyses in Appendix A.
@@ 2.3 Evaluation Metrics
where () denotes the collection of factually entailed (unentailed) sentences, and denotes the collection of answer sentences with a wrong format or without citation. While the format quality is easy to measure through heuristics, the answer’s sentence-source entailment is challenging and requires neural model prediction. In this work, we aggregate the best-performing attribution-prediction models of previous work: attrscore-flan-t5-xl and -xxl checkpoints from Yue et al. (2023) to measure entailment. To achieve higher precision, a sentence is entailed by the cited source only if both models predict “attributable”. The attributability score is not applicable for answers without any citation since models should not cite when there is no relevant source. Those answers are addressed by source quality scores (i.e., the model should cite when there is a relevant citation). We mostly follow “citation recall”, a metric introduced by Gao et al. (2023), to design attributability scores but adjust it to our stricter setting of Evidence-Based QA (more details in Appendix B).
@@ 3.2 Automated Quality Filters
The clear task definition and data creation process allow us to apply quality filters on the dataset. First, we apply a source quality filter to the original SynSciQA. Through its construction process, we know which sources are relevant and irrelevant to the question. Thus, we filter out data points that do not achieve a source quality score of one. This leaves us with 1386 samples which we call SynSciQA+.
@@ 4 Evaluation Datasets
Finally, we use another RAG tool to create ClimateQA. ClimateQA333https://huggingface.co/spaces/Ekimetrics/climate-question-answering is a RAG system that answers questions based on IPCC and IPBES reports. We pose 261 climate-related questions from Welch (2022) to the system and store the outputted sources. Again, we use this data as input for our instruction form.
@@ 5.1 Zero-Shot Performance
We first use the proposed test sets and evaluation metrics to benchmark the zero-shot performance of close-sourced and open-sourced LLMs on Evidence-Based QA. The results are shown in Table 1. We find that there is a significant performance gap between open- and close-sourced LLMs on Evidence-Based QA, although they achieve comparable performance on general instruction-following benchmarks (for instance, Zephyr-7b- vs. GPT-3.5 on MT-Bench (Tunstall et al., 2023)). Similarly, Tunstall et al. (2023) shows that Zephyr-7b- outperforms Llama-2-70b-chat on all dimensions of MT-Bench, while our evaluation shows that Llama-2-13b-chat hallucinated less than Zephyr-7b- on Evidence-Based QA. Therefore, the proposed Evidence-Based QA benchmarks can be an effective resource to benchmark LLMs’ faithfulness, supplementing MT-Bench.
@@ 5.2 SynSciQA Fine-Tuning
RQ2.2: Synthetic data as validation set for OOD performance. We observe a fluctuating performance corresponding to fine-tuning epochs. Therefore, it is important to conduct checkpoint selection over epochs with a validation (or development) set during fine-tuning. However, performance on SynSciQA is much higher than that of other OOD test sets. Therefore, the performance on a synthetic dataset cannot directly reflect the OOD or real-world performance. But can in-domain synthetic data still be an effective development set indicating which epoch may perform best on OOD data? To answer this question, we compute the Pearson’s Correlation between performance scores of different test sets. Results are presented in Table 3, illustrating that the performance on synthetic data has a strong correlation with OOD performance. However, the correlation becomes weaker when the distribution is more distant (ClimateQA and ChatReport has a weaker correlation than GenSearch). Therefore, we conclude that synthetic data can provide a valid development set for OOD performance.
@@
RQ2.3: Overfitting does exist. We suspect that fine-tuning too many epochs may cause overfitting and reduce generalisability. So we compute the Pearson Correlation between the performance scores and epoch numbers of all settings, where positive correlations indicate benefit from more epochs and negative correlations indicate overfits. Results are visualized in Figure 8, showing an overfitting trend for the majority of settings. Therefore, fine-tuning too many epochs may lead to a suboptimal performance. But we do not observe the fine-tuning overfits more to SynSciQA than others. We attribute this to SynSciQA containing different scientific topics from the training data.
@@ 5.3 Validating Attributability Score
Then, we calculate the Pearson Correlations between NLI-model-based attributability scores and human / GPT-4 evaluated attributability. As Table 4 shows, the results substantiate the validity of our method for calculating answer attributability. Correlations exceeding 80% across all comparisons between our scores and those annotated by humans and GPT-4 affirm the mutual reinforcement of the outcomes. For more details, see Appendix J.
We also notice a potential shortcut to improve the attributability score: a model may only improve its format quality (i.e., providing a citation to more sentences or more correctly writing source names) without improving the answer entailment rate. In Appendix K, we provide the entailment ratio of format-correct citations as a side result to show that this shortcut does not exist. Both format and attributability dimensions are improved by fine-tuning.
| 0.821** | 0.917** | 0.871** |
@@ 6 Related Work
Basing Answers on Sources: Prompting LLMs to respond with citations has been a popular pattern of Retrieval Augmented Generation (RAG) for better traceability (Karpukhin et al., 2020; Lewis et al., 2021; Borgeaud et al., 2021; Vaghefi et al., 2023; Ni et al., 2023; Asai et al., 2023; Li et al., 2023; Saad-Falcon et al., 2023; Gao et al., 2024). However, previous work shows that asking for citations does not make the answer more factually trustworthy (Min et al., 2023). Commercial search engines and SOTA closed-sourced LLMs suffer from unsatisfactory performance (Ni et al., 2023; Liu et al., 2023a), while open-sourced LLMs have even worse faithfulness (Yue et al., 2023; Gao et al., 2023). Therefore, LLMs, especially open-sourced ones, need essential improvement to achieve more trustworthy RAG applications.
@@ 7 Conclusion
Second, due to these budget and time limitations, we also conduct random sampling when performing human and GPT-4 evaluation to verify the attributability score instead of evaluating all instruction-answer pairs in all settings and epochs. However, we argue that the performance of different settings is uniform across different instructions and sampled data points are representative enough. Furthermore, we make all hand evaluations and generations of all checkpoints publicly available (for more details, see Appendix J).
@@
Data Privacy or Bias: There are no data privacy issues or bias against certain demographics with regard to the data collected from real-world applications and LLM generations. All artifacts we use are under creative common licence. We also notice no ethical risks associate with this work
@@ Appendix D Hand-Evaluation of the Quality Filters
On the other hand, 4% of the cases were split decisions. These predominately originate from different interpretations of nuances in the used language. Disagreements are resolved through debating about specific meanings of nuances until a concensus is achieved. Figure 10 shows an example of this.
@@ Appendix F Test Data Examples
The first evident difference lies in the statistical properties of the sources in the instructions (see Table 10). We create SynSciQA and GenSearch with the data creation pipeline described in Section 2.2. This means, we have known relevant and irrelevant sources in the datasets. On the other hand, we use top-10 retrieved sources for ChatReport and the output sources by ClimateQA for ClimateQA. As Table 10 shows, the different sourcing mechanisms result in a large difference in average length and standard deviation.
@@
Appendix G Hyperparameter and Other Settings
@@
All experiments are conducted on two clusters, one with 4 V100 GPUs and the other with 4 A100 (80G) GPUs. 1 GPU hour is used per fine-tuning.
Appendix H Relevance Label for Real RAG
Section 4 introduces that source-relevance label is available for SynSciQA thanks to the data creation process. We also annotate source-relevance for GenSearch. However, we do not annotate that for ChatReport and ClimateQA because we find most of the retrieved top-k sources in those real RAG systems are directly or indirectly relevant since they are retrieved from a narrow domain (e.g., a sustainability report). Take the following source-question pair as an example:
Question: How resilient is the organisation’s strategy when considering different climate-related scenarios, including a 2°C target or lower scenario? How resilient is the organisation’s strategy when considering climate physical risks?
Source: … Risk Management a. Describe the organization’s processes for identifying and assessing climate-related risks. CDP C2.1 CDP C2.2 CDP C2.2a Risk Management b. Describe the organization’s processes for managing climate-related risks. CDP C2.1 CDP C2.2 Risk Management c. Describe how processes for identifying, assessing, and managing climate-related risks are integrated into the organization’s overall risk management. CDP C2.1 CDP C2.2 Metrics and Targets …
Although the source does not directly address the resilience of the company’s strategy considering climate risks, it provides information about the company’s climate-related risk management, which can be indirectly useful for the resilience considering climate-related risk. Therefore, we rely on answer attributability to evaluate the real-world RAG test sets. As long as the answers have good traceability, we assume relevant information is provided to the question.
Appendix I Statistical Significance Tests
@@
Appendix J Validating the Attributability Score
The target of this validation is to reinforce the validity of the employed methodology in the answer attributability. Generally, this investigation follows the same structure as our train set validation in Appendix D. The main difference is that we now investigate all settings, including those out-of-distribution and using open-source models to find evidence that the comparisons are valid. To investigate the decisions, we repeat the evaluation of our score with both human and GPT-4 annotation. Both evaluations follow the structure of answer attributability and are articulated through the following prompt.
@@
Appendix K Format Short-Cut in Attributability